1. Suppose country A can produce 30 units of Corn per time period (if it specializes in Corn). Country A can produce 4 outfits per time period (if it specializes in outfits). If country B specializes in Corn, it can produce 40 units per time period. If country B specialized in Outfits, it can produce 3 outfits per time period.

A has a comparative advantage in Corn.

B has a comparative advantage in Outfits.

neither country has a comparative advantage.

A has a comparative advantage in Outfits.

none of the other options.

2.With regard to the previous questions, which of the following could be the world equilibrium relative price of Outfits (O) in terms of Corn (C) (PO/PC) if trade opened up and countries A and B specialized according to comparative advantage?

Group of answer choices

5

16

10

7

14

3. In a setups like the previous questions, we have said that if both specialize in the good in which they have a comparative advantage, then gains from trade are possible. The primary reason this possibility arises in this setting is

Group of answer choices

that one country has an absolute advantage.

that one country's production possibilities frontier (PPF) is not a straight line.

that one country's PPF lies completely outside the other country's PPF.

the difference in production tradeoffs.

that neither country has an absolute advantage.

4.Country A can produce one Outfit in 12 hours and one unit of Corn in 2 hours. Country B can produce one Outfit in 16 hours and one unit of Corn in 3 hours. If specialization according to the principle of comparative advantage were to occur, we should expect

   

Group of answer choices

country A to specialize in Outfits and to import Corn.

country B to specialize in Outfits and to import Outfits.

country A to specialize in Corn and to import Outfits.

country A to specialize in Corn and to export Outfits.

country B to specialize in Outfits and to export Corn.

5.Suppose the labor endowment in country A is LA = 500. Suppose that the labor input requirement for good X is  aX A = 10. Suppose that the labor input requirement for good Y is aY A = 5. Then which of the following statements is correct?

Group of answer choices

Country A can produce at most 100 units of good Y.

Country A can produce at most 100 units of good X.

Country A must give up 5 units of Y production in order to increase X production by 1.

None of the other options.

Country A must give up 0.5 units of Y production in order to increase X production by 1.
Answer #1

Ans-1) A has comparative advantage in outfits.

working : table for comparative advantage on opportunity costs basis-:

Country CORN OUTFIT
A 30 4
B 40 3
Country

CORN OPPORTUNITY

COST

OUTFIT OPPORTUNITY

COST

A 4/30= 0.133 30/4=7.5
B 3/40= 0.075 40/3=13.33

Form the opportunity cost table , we can say that Country B has less opportunity cost of producing corn and Country A has less opportunity cost of producing outfit.

Therefore , Country A has comparative advantage in Outfit and Country B has comparative advantage in Corn .